on cathode ray oscilloscope,the distance between crests is 2cm.if the time base is set at 10ms/cm.what is the period of the wave
1
Expert's answer
2020-05-04T12:51:12-0400

The time between two peaks of electrical vibrations is calculated using a simple formula

"time=distance\cdot timebase=2cm\cdot 10ms/cm=20ms=2.0\cdot10^{-2}s"

Answer: The period of the wave is "2.0\cdot 10^{-2}s"